# A Study on Adversarial Robustness of Discriminative Prototypical Learning

### v1

@@ 3.3 Dual-Branch Training Procedure
Clean Data Branch. In this branch designated for clean inputs $x_{i}$, both the feature extractor parameters $\theta$ and the corresponding prototypes $c_{y_{i}}$ are updated during backpropagation. So, this branch adapts the prototypes to the original unperturbed input instances, ensuring well-structured class clusters in the feature space and maintaining stable representations of each class center.
@@
7:     for each mini-batch =$\mathcal{B}=\{(x_{i},y_{i})\}_{i=1}^{B}\subset\mathcal{D}$ do
@@
11:              Compute feature representations $f(\tilde{x_{i}};\theta)$ and probabilities $p_{j}(\tilde{x_{i}})$
12:              Compute the loss functions $L_{\text{DPP}}({x_{i}},{y_{i}})$, $L_{\text{DPP}}(\tilde{x_{i}},{y_{i}})$ and $L_{\text{DFA}}({x_{i}},\tilde{x_{i}})$
@@
22:              $c_{j}\leftarrow c_{j}-\eta\nabla_{c_{j}}(\lambda_{\text{DNP}}L_{\text{DNP}}+\frac{1}{2B}\sum_{i}^{B}L_{\text{DPP}}(x_{i},y_{i})+\lambda_{\text{DFA}}L_{\text{DFA}}(x_{i},\tilde{x_{i}}))$
@@ 4 Experiments
In this section, we extensively evaluate our proposed Adv-DPNP framework on two widely used datasets with adversarial robustness, namely CIFAR-10 [15] and CIFAR-100 [15]. We use the ResNet-18 architecture [1] in all experiments for these datasets. Section 4.1 explains our experimental setup, including parameter settings and training protocols. In Section 4.2, we present a comparative robustness analysis of our proposed Adv-DPNP against the most prominent adversarial defense methods, including AT [10], TRADES [11], MART [12], and AT_HE [22]. Subsequently, Section 4.3 examines the effectiveness of Adv-DPNP in structuring the feature space for robust and discriminative representations. Section 4.4 explores the robustness of our framework in common image corruptions. Finally, in Section 4.5, we discuss potential gradient obfuscation issues and show how our method addresses them.
@@ 4.1 Experimental Setup
Training Parameters. We follow the recommended settings from [37] for adversarially training the models to prevent the training parameters from undermining robustness improvements. Specifically, all models are trained on ResNet-18 [1] with a weight decay of $5\times 10^{-4}$ for 110 epochs. We use Stochastic Gradient Descent (SGD) with a momentum of 0.9 and an initial learning rate of 0.1, which is reduced to 0.01 at epoch 100 and 0.001 at epoch 105. The batch size is set to 128 in all experiments. We apply standard data augmentation techniques such as random cropping and horizontal flipping.
Adversarial Training Setup. For adversarial training, we generate adversarial examples using the PGD attack [10] with the $l_{\infty}$-norm bound $\epsilon=8/255$, step size $2/255$, and 10 iterations. These settings are consistent with standard practice for CIFAR-10 and CIFAR-100, aiming to provide a fair comparison with other adversarial defense methods.
Hyperparameters. For baseline methods, we use the hyperparameters suggested in their respective papers: $61/\lambda=$ 6 for TRADES, $\lambda=5$ for MART, and $s=15$, $m=0.2$ for AT_HE. Regarding our Adv-DPNP framework, we set $\alpha=40$ for both CIFAR-10 and CIFAR-100. After a small empirical search, we found that on CIFAR-10, the values $\lambda_{\text{DPP}}=0.1$, $\lambda_{\text{DNP}}=0.1$, and $\lambda_{\text{DFA}}=2$ perform reliably, while on CIFAR-100, setting $\lambda_{\text{DPP}}=0.3$, $\lambda_{\text{DNP}}=3$, and $\lambda_{\text{DFA}}=2.4$ yields stable performance. In the next sections, we report detailed comparisons under these configurations and demonstrate the effectiveness of Adv-DPNP in achieving robust and discriminative feature representations.
@@ 4.2 Robustness Against Adversarial Perturbations
Evaluation Details. For all baselines and datasets, we report the accuracy of each model on clean (natural) samples and under adversarial attacks. Specifically, we evaluate performance using $l_{\infty}$-constrained PGD attacks ($\epsilon=8/255$, 20 steps with a step size of $1/255$), $l_{2}$-constrained PGD attacks ($\epsilon=128/255$, 20 steps with a step size of $16/255$), and $l_{1}$-constrained PGD attacks ($\epsilon=2000/255$, 100 steps with a step size of 50). Furthermore, we assess the performance of trained models using AA [21], an ensemble of white-box and black-box attacks, to provide a comprehensive and reliable evaluation of adversarial robustness across various models.
@@
Table 1: Comparison of clean and adversarial accuracy (%) on CIFAR-10 under various PGD attacks ($l_{1}$,$l_{2}$,$l_{\infty}$) and AutoAttack (AA). The ”Avg” column denotes the average of clean and AA accuracy. Our Adv-DPNP achieves superior performance and offers a better balance between clean accuracy and robustness compared to other baseline methods.
Clean PGD AA Avg
- $l_{1}$ $l_{2}$ $l_{\infty}$ $l_{\infty}$
ST 94.39 40.77 0.04 0.0 0.0 47.19
AT 83.78 51.04 63.72 52.75 48.20 65.99
AT_HE 80.28 48.44 60.67 50.77 47.29 63.78
TRADES 81.38 49.72 62.37 53.18 49.00 65.19
MART 79.60 52.39 61.41 51.94 48.18 63.89
Adv-DPNP 84.63 55.11 64.13 53.68 49.18 66.90
[…]
@@
Performance. We compare our proposed Adv-DPNP with the state-of-the-art adversarial defense methods on CIFAR-10 and CIFAR-100 as shown in Tables 1 and 2. Several key observations can be drawn from these results. First, Adv-DPNP noticeably increases clean accuracy for both datasets, improving by 0.85% on CIFAR-10 and 1.78% on CIFAR-100 compared to other defenses. This improvement indicates that the dual-branch training scheme effectively captures true class prototypes from clean data, leading to more accurate predictions in natural settings. Second, our method outperforms the baseline approaches in unforeseen perturbation types in both datasets. In particular, Adv-DPNP boosts robustness against $l_{1}$ perturbations by 2.7% for CIFAR-10 and robustness against $l_{2}$ perturbations by 1.07% for CIFAR-100, suggesting that our proposed Adv-DPNP has a better transferability of robustness between $l_{\infty}$ to $l_{1}$ and $l_{2}$ perturbations. Third, when considering the average of clean and AA accuracy indicated by the ”Avg” column in Tables 1 and 2, Adv-DPNP achieves a better balance between clean accuracy and robustness. Specifically, Adv-DPNP obtains the highest AA accuracy on CIFAR-10 and improves the average accuracy by 0.91%, while on CIFAR-100 it enhances the average accuracy by 1.05% and maintains competitive AA performance.
Table 2: Comparison of clean and adversarial accuracy (%) on CIFAR-100 under multiple PGD attacks ($l_{1}$,$l_{2}$,$l_{\infty}$) and AutoAttack (AA). The “Avg” column denotes the average of clean and AA accuracy. Despite the increased difficulty of CIFAR-100, our Adv-DPNP consistently outperforms other adversarial defenses in terms of both robustness and natural accuracy.
Clean PGD AA Avg
- $l_{1}$ $l_{2}$ $l_{\infty}$ $l_{\infty}$
ST 76.24 25.90 01.42 00.03 00.00 38.12
AT 57.91 38.03 38.95 29.17 24.98 41.45
AT_HE 57.11 36.71 37.99 29.35 25.73 41.42
TRADES 57.99 38.43 38.77 30.17 25.06 41.53
MART 53.64 34.38 37.66 29.46 25.61 39.63
Adv-DPNP 59.77 39.26 40.02 30.80 25.39 42.58
[…]
@@
Robustness against different perturbation sizes. To further evaluate the robustness of adversarially trained models in Table 1, we report their accuracy against $l_{\infty}$ constraint PGD and AA with different perturbation sizes. As shown in Table 3, our proposed Adv-DPNP consistently outperforms baseline methods across all tested perturbation sizes, further confirming its robust improvement.
These findings demonstrate that aligning adversarial features with prototypes derived from clean data can effectively narrow the gap between clean accuracy and adversarial robustness.
Table 3: Adversarial accuracy (%) under different $l_{\infty}$ perturbation bounds (ϵitalic-$\epsilon\in$2/255,4/255,6/255) evaluated via PGD and AutoAttack (AA) on CIFAR-10. Adv-DPNP maintains higher robustness across all tested $\epsilon$ values.
$\epsilon=\frac{2}{255}$ $\epsilon=\frac{4}{255}$ $\epsilon=\frac{6}{255}$
@@
AT_DPNP 78.07 77.25 70.71 68.76 62.13 58.85
[…]
4.3 Analysis of Feature Space Organization
To investigate how our Adv-DPNP framework preserves discriminative structures under adversarial conditions, we analyze the learned feature spaces of adversarially trained models explained in Section 4.1. Specifically, we use four complementary metrics including Fisher’s Discriminant Ratio (FDR) [38], Angular Fisher Score (AFS) [39], Separation to Compactness Ratio (SCR) [14], and Class Center Angular Separation [14] to evaluate feature discrimination in both clean and adversarial scenarios. These metrics collectively capture different aspects of intra-class compactness and inter-class separation, providing a comprehensive perspective on each model’s ability to maintain a robust and well-organized latent space.
@@ 4.3 Analysis of Feature Space Organization
where $tr(.)$ is the trace operation, $S_{j}^{w}$ and $S_{j}^{b}$ denote the within-class and between-class scatter matrices for class $j$ respectively, $\mu_{j}=\frac{1}{N_{j}}\sum_{x_{i}\in\mathcal{C}_{j}}\,f(x_{i};\theta)$ is the mean feature vector of class $j$, and $\mu=\tfrac{1}{N}\sum_{i=1}^{N}\,f(x_{i};\theta)$ is the global mean feature vector. A lower FDR value indicates a more discriminative feature space, with smaller within-class variance and larger between-class variance.
AFS. In the most DNNs trained with softmax function, classification is often based on the angular similarity between the feature vectors and the classifier weights of each class. This motivates the use of AFS to quantitatively assess the feature discrimination in angular space by computing the ratio of angular within-class scatter to angular between-class scatter:
@@
SCR. Unlike FDR and AFS which measure global variance, SCR explicitly considers the distance between each class center and its nearest rival. Then, SCR is given by:
@@
where $c_{j}$ denote the center for class $j$. For baseline methods, the classifier weights are used as class centers. A higher SCR indicates that each class is better separated from its nearest rival while maintaining intra-class compactness.
@@
Both metrics are computed in degrees. Higher values of MeanSep and MinSep indicate better separation between class centers, which enhances margin-based discrimination.
Table 4: Quantitative evaluation of feature-space discrimination across different methods on CIFAR-10. The models are those evaluated in Section 4.2. Arrows ($\uparrow$ / $\downarrow$) indicate whether higher or lower values are better. The best result in each column is in bold, and the second-best is underlined.
[…]
cln adv cln adv cln adv - -
ST 0.029 0.597 0.200 2.850 0.577 0.252 82.24 75.52
@@
[…]
Table 5: Quantitative evaluation of feature-space discrimination across different methods on CIFAR-100. The models are those evaluated in Section 4.2. Arrows ($\uparrow$ / $\downarrow$) indicate whether higher or lower values are better. The best result in each column is in bold, and the second-best is underlined.
[…]
cln adv cln adv cln adv - -
ST 0.011 0.114 0.389 1.479 0.190 0.066 65.95 44.96
@@
[…]
Results. We evaluate these metrics for both clean and adversarial features, where adversarial examples are generated using the 20-step $l_{\infty}$-PGD attack. The results for CIFAR-10 and CIFAR-100 are reported in Tables 4 and 5, respectively. Across all models, adversarial features exhibit higher FDR and AFS values compared to clean features, as adversarial perturbations tend to shift the feature representations towards incorrect class centers [33]. Although the undefended model (ST) achieves the lowest FDR and AFS values for clean features, it fails to maintain a consistent latent space structure under adversarial perturbations. In contrast, adversarially trained models preserve better structure when confronted with adversarial examples. Notably, Adv-DPNP exhibits the lowest FDR and AFS values among robust models in both clean and adversarial settings, indicating that Adv-DPNP generates angularly discriminant features and reduces radial variance. Additionally, Adv-DPNP maintains the highest SCR and class separation metrics (MeanSep and MinSep), which demonstrates better separability compared to other methods. These results show that Adv-DPNP successfully preserves a discriminative latent space, exhibiting high intra-class compactness and inter-class separation in both clean and adversarial settings.
@@
To further illustrate the effectiveness of Adv-DPNP in preserving a discriminative and robust latent space, we conduct an additional experiment on the MNIST dataset [40]. We employ the LeNets++ architecture [30], utilizing its two-dimensional feature representation for direct visualization. Specifically, we train AT as a baseline model and our proposed Adv-DPNP. The resulting two-dimensional distributions of clean and adversarial features from the test dataset are visualized in Figure 2. Compared to standard AT, Adv-DPNP significantly reduces both angular and radial variations of feature representations under clean and adversarial conditions. Also, Adv-DPNP creates more distinct inter-class margins, yielding improved separation between class prototypes. Moreover, the adversarially perturbed features in Adv-DPNP closely align with the prototypes learned from clean data, demonstrating its capability to robustly preserve the geometric structure of the feature space.
4.4 Robustness on Common Corruptions
To assess the natural robustness of our method, we evaluate the performance of adversarially trained models on the CIFAR-10 common corruptions dataset (CIFAR-10-C) [41]. This dataset is created by corrupting the original CIFAR-10 images with 15 test corruptions and 4 validation corruptions. These corruptions are categorized into four types including noise, blur, weather, and digital, with each type having five levels of severity or intensity.

### v2

@@ 3.3 Dual-Branch Training Procedure
Clean Data Branch. In this branch designated for clean inputs $x_{i}$, both the feature extractor parameters $\theta$ and the corresponding prototypes $c_{y_{i}}$ are updated during backpropagation. So, this branch adapts the prototypes to clean inputs, ensuring well-structured class clusters in the feature space and maintaining stable representations of each class center.
@@
7:  for each mini-batch $\mathcal{B}=\{(x_{i},y_{i})\}_{i=1}^{B}\subset\mathcal{D}$ do
@@
[…]
@@
22:     $)c_{j}\leftarrow c_{j}-\eta\nabla_{c_{j}}(\lambda_{\text{DNP}}L_{\text{DNP}}+\frac{1}{2B}\sum_{i}^{B}L_{\text{DPP}}(x_{i},y_{i})+\lambda_{\text{DFA}}L_{\text{DFA}}(x_{i},\tilde{x_{i}}$))
@@ 4 Experiments
In this section, we extensively evaluate Adv-DPNP on three widely used datasets, namely CIFAR-10 [24], CIFAR-100, and SVHN [35]. For a comprehensive comparison, we conduct experiments using two network backbones: ResNet-18 (RN-18) and WideResNet-34-10 (WRN-34-10). Section 4.1 explains our experimental setup, including parameter settings and training protocols. In Section 4.2, we present a comparative robustness analysis of Adv-DPNP against several prominent baseline methods, including AT [30], TRADES [52], MART [42], and AT_HE [36], as well as SOTA methods such as OC [46], PCL [33], FSR [21], HAT [38], DIO [10], and IKL [6]. Section 4.3 analyzes the effectiveness of Adv-DPNP in structuring the learned feature space for discriminative representations. Section 4.4 examines the robustness under common image corruptions. Finally, Section 4.5 presents experiments designed to rule out the possibility of gradient obfuscation in Adv-DPNP.
@@ 4.1 Experimental Setup
Training Parameters. We follow the recommended settings from [37] for adversarially training the models to prevent the training parameters from undermining robustness improvements. Specifically, all evaluated models are trained with a weight decay of $5\times 10^{-4}$ for 110 epochs. We use stochastic gradient descent with a momentum of 0.9 and an initial learning rate of 0.1, which is reduced to 0.01 at epoch 100 and 0.001 at epoch 105. The batch size is set to 128 in all experiments. We also apply standard data augmentation during training, namely random cropping and horizontal flipping. All models, including Adv-DPNP and all baseline/SOTA defenses, are trained from scratch under this training protocol, using the authors’ official implementations to ensure a fair comparison.
Adversarial Training Setup. For adversarially training the models, we follow standard practice and generate adversarial examples using the PGD attack [30] under the $l_{\infty}$ norm bound $\epsilon=8/255$ for 10 iterations. The step size is set to $2/255$ for CIFAR-10/100 and $1/255$ for SVHN. This setup is used for all adversarially trained methods to ensure comparability.
Hyperparameters. For all evaluated methods, we use the hyperparameters suggested in their respective papers. For Adv-DPNP, we set the hyperspherical radius to $\alpha=40$ for all experiments. On CIFAR-10 and SVHN, we use $\lambda_{\text{DPP}}=0.1$, $\lambda_{\text{DNP}}=0.1$, and $\lambda_{\text{DFA}}=2$, while on CIFAR-100, we set $\lambda_{\text{DPP}}=0.3$, $\lambda_{\text{DNP}}=3$, and $\lambda_{\text{DFA}}=2.4$ to account for its larger number of classes and tighter class separations. These values are kept constant across all runs and comparisons.
@@ 4.2 Robustness Against Adversarial Perturbations
Evaluation Details. We evaluate the adversarial robustness of Adv-DPNP using a variety of widely adopted attacks. To thoroughly assess the robustness from multiple perspectives, we perform comparisons across two complementary sets of rival methods. The first set, which we refer to as baselines, consists of famous defense methods which are tested under the standard PGD attacks constrained by $l_{\infty}$ ($\epsilon=8/255$, with a step size of $1/255$ for 20 iterations), $l_{2}$ ($\epsilon=128/255$ with a step size of $16/255$ for 20 iterations), and $l_{1}$ ($\epsilon=2000/255$ with a step size of 50 for 100 iterations) norms, along with AutoAttack (AA), which is an ensemble of three white-box attacks and one black-box attack. We further evaluate our Adv-DPNP against the second set which includes recently published SOTA approaches under widely used diverse adversarial attacks, including FGSM, PGD (20 and 100 iterations), C&W (30 iterations with a confidence margin of 50), and an Ensemble attack. The Ensemble accuracy is calculated by sequentially applying FGSM, PGD-20, PGD-100, and C&W attacks to each test sample, marking it as failed if its predicted label changes under any of these attacks.
CIFAR‑10 CIFAR‑100
Method Clean AA Avg PGD Clean AA Avg PGD
$l_{\infty}$ $l_{1}$ $_{2} ll_{\infty}$ $l_{\infty}$ $l_{1}$ $_{2} ll_{\infty}$
ST 94.39 0.00 47.20 40.77 0.04 0.00 76.24 0.00 38.12 25.90 1.42 0.03
AT 83.78 48.20 65.99 51.04 63.72 52.75 57.91 24.98 41.45 38.03 38.95 29.17
AT_HE 80.28 47.29 63.79 48.44 60.67 50.77 57.11 25.73 41.42 36.71 37.99 29.35
TRADES 81.38 49.00 65.19 49.72 62.37 53.18 57.99 25.06 41.53 38.43 38.77 30.17
MART 79.60 48.18 63.89 52.39 61.41 51.94 53.64 25.61 39.63 34.38 37.66 29.46
Adv‑DPNP 84.63 49.18 66.91 55.11 64.13 53.68 59.77 25.39 42.58 39.26 40.02 30.80
@@
Table 1: Comparison of clean and adversarial accuracy (%) on CIFAR‑10 and CIFAR‑100 under AutoAttack (AA) and various PGD attacks ($l_{1}$$_{2},ll_{\infty}$). The “Avg” column denotes the average of clean and AA accuracy. Adv-DPNP offers a better balance between clean accuracy and robustness compared to other baseline methods.
@@
Comparison with Baselines. In Table 1, we compare the proposed Adv-DPNP with the baseline adversarial defenses on CIFAR-10 and CIFAR-100. For reference, the standard training (ST) model (i.e., an undefended baseline) achieves the highest clean accuracy but collapses under attacks (AA $\approx$ 0%); thus, we compare clean accuracy only among adversarial defenses. As shown in Table 1, Adv-DPNP increases clean accuracy on both datasets, improving by 0.85% on CIFAR-10 and 1.78% on CIFAR-100 compared with other defenses. This improvement suggests that the dual-branch training scheme in Adv-DPNP effectively captures the true class prototypes from clean data. Furthermore, Adv-DPNP outperforms the baselines under unforeseen perturbation types on both datasets. In particular, it improves robustness against $l_{1}$ perturbations by 2.7% on CIFAR-10 and against $l_{2}$ perturbations by 1.07% on CIFAR-100, indicating better transferability of robustness from $l_{\infty}$ to $l_{1}$ and $l_{2}$ norms. Lastly, considering the average of clean and AA accuracy (denoted by “Avg” in Table 1), Adv-DPNP achieves a better balance between clean accuracy and robustness. Specifically, it attains the highest AA accuracy on CIFAR-10 and improves Avg by 0.92%, while on CIFAR-100 it increases Avg by 1.05% and maintains competitive AA performance.
Dataset / Backbone Method Clean AA Avg FGSM PGD-20 PGD-100 C&W Ensemble
Accuracy (%)
CIFAR-10 / RN-18 OC 81.80 46.46 64.13 57.14 51.40 49.99 50.28 49.59
PCL 83.70 40.85 62.28 51.23 44.63 43.21 43.69 41.90
HAT 84.28 47.51 65.90 57.85 52.01 50.80 50.21 48.68
FSR 81.83 47.33 64.58 57.03 52.52 51.28 49.93 48.38
DIO 83.18 48.21 65.70 58.06 52.59 51.25 50.90 49.25
IKL 77.90 48.88 63.39 57.07 54.14 53.65 49.93 49.61
Adv-DPNP 84.63 49.18 66.91 58.12 53.68 52.61 50.97 50.14
CIFAR-100 / RN-18 OC 59.24 24.49 41.87 33.02 27.36 26.10 26.43 25.81
PCL 59.08 20.61 39.85 30.64 26.75 26.07 23.77 22.18
HAT 59.60 22.79 41.20 29.61 26.35 25.44 25.05 23.81
FSR 58.04 23.38 40.71 30.64 27.24 26.40 25.72 24.19
DIO 54.39 24.50 39.45 31.83 29.35 28.79 26.63 25.68
IKL 56.62 25.43 41.03 32.84 31.17 30.70 26.42 26.29
Adv-DPNP 59.77 25.39 42.58 33.19 30.80 30.15 27.29 26.58
SVHN / RN-18 FSR 92.96 45.22 69.09 65.13 52.82 49.43 48.97 47.05
DIO 93.52 45.44 69.48 66.17 52.97 49.48 49.72 47.57
IKL 85.05 47.53 66.29 61.46 55.50 54.46 50.73 50.36
Adv-DPNP 93.06 48.04 70.55 68.43 57.45 54.85 51.54 50.53
CIFAR-10 / WRN-34-10 FSR 86.39 51.14 68.77 61.69 55.76 54.28 54.19 52.37
DIO 85.28 51.95 68.62 60.68 54.75 53.41 54.79 52.70
IKL 82.23 53.10 67.67 61.50 57.73 57.26 54.36 53.87
Adv-DPNP 87.47 52.77 70.12 62.14 57.26 55.07 54.96 53.74
Table 2: Robustness comparison on three datasets and two backbones. Columns list accuracy (%) on clean data and six diverse attacks. The Ensemble column counts a sample as wrong if any of FGSM, PGD-20, PGD-100 or C&W succeeds, providing a fast worst-case estimate. “Avg” is the mean of Clean and AutoAttack (AA) accuracies. Adv-DPNP attains the highest “Avg” in all four settings and never drops below second place in AA, demonstrating consistent gains over the competing methods.
Comparison with SOTA. Table 2 further compares Adv-DPNP against the SOTA defenses across three datasets (CIFAR-10, CIFAR-100, SVHN) and two backbone architectures (ResNet-18 and WRN-34-10). Several important observations emerge from this evaluation. First, Adv-DPNP consistently achieves the highest Avg in all four evaluation blocks. Specifically, on CIFAR-10 with ResNet-18, it achieves an Avg of 66.91%, outperforming the second-best margin-based defense (HAT) by 1.01%. On CIFAR-100 with ResNet-18, it attains an Avg of 42.58%, surpassing the next-best method by 0.72%. On SVHN with ResNet-18, Adv-DPNP again improves the Avg to 70.55%, outperforming the orthogonality-based defense DIO by 1.07%. Finally, with WRN-34-10 trained on CIFAR-10, it achieves an Avg of 70.12%, surpassing the feature-masking method (FSR) by 1.35%. Second, Adv-DPNP demonstrates competitive robustness under AA, consistently achieving either the highest or second-highest AA accuracy across all four settings. This indicates that the robustness gains generalize across datasets and architectures. Third, Adv-DPNP consistently ranks among the top defenses against diverse individual attacks (FGSM, PGD-20, PGD-100, C&W) as well as under the Ensemble metric. Notably, it achieves the highest Ensemble accuracy in three of the four evaluation settings.
@@
Robustness against different perturbation sizes. To further evaluate the robustness of adversarially trained models, we report their accuracy against $l_{\infty}$ constrained PGD and AA with different perturbation sizes. As shown in Table 3, Adv-DPNP consistently outperforms other methods across all tested perturbation sizes, further confirming its robust improvement.
Method $\epsilon=\frac{2}{255}$ $\epsilon=\frac{4}{255}$ $\epsilon=\frac{6}{255}$
@@
FSR 75.70 74.40 68.51 65.82 60.62 56.90
DIO 77.28 76.37 70.04 68.11 61.78 58.72
IKL 72.74 71.74 66.83 64.60 60.69 57.18
Adv_DPNP 78.07 77.25 70.71 68.76 62.13 58.85
Table 3: Adversarial accuracy (%) evaluated via PGD and AutoAttack (AA) under different $l_{\infty}$ perturbation bounds ($\epsilon\in\{2/255,4/255,6/255\}$) on CIFAR-10. Adv-DPNP maintains higher robustness across all tested $\epsilon$ values.
Collectively, these observations confirm the effectiveness and broad generalizability of the Adv-DPNP framework. By anchoring adversarially perturbed representations to prototypes learned solely from clean data, Adv-DPNP successfully shrinks the gap between accuracy and robustness, exhibiting robust performance across multiple datasets, attacks, and network architectures.
4.3 Discriminative Aspects of the Feature Space
To investigate how Adv-DPNP preserves discriminative feature representations under adversarial perturbations, we analyze the learned feature spaces of adversarially trained models explained in Section 4.1, comparing Adv-DPNP against conventional defenses. Specifically, we use four complementary metrics including Fisher’s Discriminant Ratio (FDR) [12], Angular Fisher Score (AFS) [28], Separation to Compactness Ratio (SCR) [49], and Class Center Angular Separation [49] to evaluate feature discrimination in both clean and adversarial scenarios. These metrics collectively capture different aspects of intra-class compactness and inter-class separation, providing a comprehensive perspective on each model’s ability to maintain a robust and well-organized latent space.
@@ 4.3 Discriminative Aspects of the Feature Space
where $tr(.)$ is the trace operation, $S_{j}^{w}$ and $S_{j}^{b}$ denote within-class and between-class scatter matrices for class $j$ respectively, $\mu_{j}=\frac{1}{N_{j}}\sum_{x_{i}\in\mathcal{C}_{j}}\,f(x_{i};\theta)$ is the mean feature vector of class $j$, and $\mu=\tfrac{1}{N}\sum_{i=1}^{N}\,f(x_{i};\theta)$ is the global mean feature vector. A lower FDR value indicates a more discriminative feature space, with smaller within-class variance and larger between-class variance.
AFS. In most DNNs trained with the softmax function, classification is based on the angular similarity between the feature vectors and the classifier weights of each class. This motivates the use of AFS to quantitatively assess the feature discrimination in angular space by computing the ratio of angular within-class scatter to angular between-class scatter:
@@
SCR. Unlike FDR and AFS, which measure global variance, SCR explicitly considers the distance between each class center and its nearest rival. Then, SCR is given by:
@@
where $c_{j}$ denotes the center for class $j$. For competing methods, classifier weights are used as class centers. A higher SCR indicates that each class is better separated from its nearest rival while maintaining intra-class compactness.
@@
Both metrics are computed in degrees. Larger MeanSep and MinSep indicate better separation between class centers, implying stronger margin-based discrimination.
Dataset Method FDR $\downarrow$ AFS $\downarrow$ SCR $\uparrow$ MeanSep $\uparrow$ MinSep $\uparrow$
cln adv cln adv cln adv
CIFAR-10 ST 0.029 0.597 0.200 2.850 0.577 0.252 82.24 75.52
@@
CIFAR-100 ST 0.011 0.114 0.389 1.479 0.190 0.066 65.95 44.96
@@
Table 4: Quantitative evaluation of feature-space discrimination across different methods on CIFAR-10 and CIFAR-100. Arrows ($\uparrow$ / $\downarrow$) indicate whether higher or lower values are better. The best result in each column is in bold, and the second-best is underlined.
Results. We evaluate these metrics on both clean and adversarial features, where adversarial examples are generated using the 20-step $l_{\infty}$-PGD attack. The results for CIFAR-10 and CIFAR-100 are reported in Table 4. Across all models, adversarial features exhibit higher FDR and AFS values compared to clean features, as adversarial perturbations tend to shift the feature representations towards incorrect class centers [31]. As expected, undefended ST achieves the lowest FDR and AFS values for clean features, but it drastically fails to maintain a consistent latent space structure under adversarial perturbations, leading to poor robustness and geometric stability. In contrast, adversarially trained models preserve better structure when confronted with adversarial examples. Notably, our Adv-DPNP exhibits the lowest FDR and AFS values among robust models in both clean and adversarial settings, indicating better Euclidean and angular discrimination. Additionally, Adv-DPNP maintains the highest SCR and the best angular margins (MeanSep/MinSep reach $95.32^{\circ}/94.51^{\circ}$ on CIFAR-10 and $89.90^{\circ}/89.52^{\circ}$ on CIFAR-100, substantially better than other methods), increasing its generalization. These results indicate that Adv-DPNP successfully learns a globally well-separated set of anchors and preserves a discriminative latent space, exhibiting high intra-class compactness and inter-class separation in both clean and adversarial settings. Moreover, the smaller change between corresponding values of clean/adversarial columns in the case of Adv-DPNP suggests that our dual-branch training procedure successfully learns a robust feature extractor and discriminant prototypes.
@@
To further illustrate the discrimination in the latent space, we conduct an additional experiment on the MNIST dataset [26]. We employ the LeNets++ architecture [44], utilizing its two-dimensional feature representation for direct visualization. Specifically, we train AT as a baseline model and our proposed Adv-DPNP. The resulting two-dimensional distributions of clean and adversarial features from the test dataset are visualized in Figure 2. Compared to standard AT, Adv-DPNP significantly reduces both angular and radial variations of feature representations under clean and adversarial conditions. Also, Adv-DPNP creates more distinct inter-class margins, yielding improved separation between class prototypes. Moreover, the adversarially perturbed features in Adv-DPNP closely align with the prototypes learned from clean data, demonstrating its capability to robustly preserve the geometric structure of the feature space.
4.4 Robustness against Common Corruptions
Figure 3: Comparison of models’ performance against five different corruption severity levels for the CIFAR-10-C dataset. Each bar represents an average of 19 corruption types for a given level of corruption strength. The Y-axis is limited to 60–85% for readability.
Method AT AT_HE TRADES MART PCL OC HAT FSR DIO IKL Adv-DPNP
Avg Acc (%) 76.16 73.20 73.84 69.04 75.54 74.09 75.92 74.06 75.64 69.40 76.73